# Understanding Ice Crystal Habit Diversity with Self-Supervised Learning

### v1

@@ 1 Introduction
Clouds are one of the largest sources of uncertainty in climate models (Morrison et al., 2020; Lamb et al., 2025). They are notoriously difficult to represent accurately in models, and ice-containing clouds are especially challenging due to highly diverse properties such as crystal morphology (Järvinen et al., 2018). Ice microphysical properties alter particle-radiation interactions and aerodynamics at the single-particle scale; and influence global radiative forcing, precipitation, and spatiotemporal distributions of clouds through a cascade of multiscale interactions (Shima et al., 2020; Chandrakar et al., 2024). Improving our understanding of clouds is crucial, since uncertainties in future cloud behavior largely drive the overall uncertainty of future climate projections (Bony et al., 2006; Sherwood et al., 2020; Tan et al., 2025).
One important way to constrain ice microphysical properties is to take in situ measurements. For example, to understand the distribution of ice crystal habit (i.e., shape), millions of images of cloud particles have been taken on numerous airborne campaigns using cloud particle imagers (CPI) (Baker and Lawson, 2006). Historically, image processing techniques have been used to extract microphysically-relevant properties from CPI images (McFarquhar et al., 2017; Lawson et al., 2019), and more recently, supervised ML has been used to improve predictions of particle properties (Przybylo et al., 2022). However, unsupervised ML has largely been underutilized in the context of analyzing CPI data and in situ microphysical observations more broadly.
@@ 2.2 Efficient self-supervised pre-training
Self-supervised learning (SSL) is an effective approach to learn informative representations of unlabeled data (Ericsson et al., 2022; Kolesnikov et al., 2019). Such representations enable various downstream analyses as well as efficient predictions with minimum labeling efforts (Assran et al., 2022). CPI images exhibit natural clusters characterized by ice habits. Hence, we consider state-of-the-art, clustering-based SSL methods from the DINO family (Caron et al., 2021; Zhou et al., 2022), which are trained to assign each image to a cluster such that its augmented views, obtained by applying data augmentations, are also assigned to the same cluster. Govindarajan et al. (2023) showed that the representations constitute a mixture model of von Mises-Fisher (vMF) distributions. We use the iBOT-vMF method (Govindarajan et al., 2023) to pre-train our models and use the small Vision Transformer model architecture with a patch size of $16$ (more details in A.5.1). We evaluate the learned representations with the downstream task of classifying the CPI-21K dataset.
@@
The above experiments showed that data curation is important and that ImageNet pre-trained models work reasonably well on CPI data. Also, pre-training on well-curated datasets in the target domain demonstrated potential for performance improvements with better downstream classification results. With this motivation, we investigated if we could pre-train models for CPI data more efficiently. Specifically, we pre-train a model on the curated CPI-H-1M dataset using iBOT-vMF for only 10 epochs and initialize the model with the ImageNet pre-trained model weights. Given a well-curated dataset like CPI-H-1M, this approach is $\sim 30\times$ more compute efficient than directly pre-training on the large CPI-3M dataset and resulted in the best performance based on validation of learned representations (see Table 1).
Table 1: Comparison of self-supervised learning models on the task of classifying the CPI-21K dataset using kNN and logistic regression classifiers. We report the Top-1 accuracy metric.
@@ 3.1 Quality of learned representations
To supplement the validation results in Table 1, here we (1) confirm if clusters match expert habit labels, and (2) compare to a feature-extraction based baseline to demonstrate the benefit of SSL representations. For (1), we used dimensionality reduction to inspect clusters in 2D space. For clarity, we used a balanced subset of 3000 CPI-3M samples, and three of the seven classes to reduce visual clutter (see Appendix A.1). UMAP and PCA were used to project the 384-dimensional SSL embeddings into 2D. Figure 2 visualizes the projections, with points colored by labels predicted by a CNN from Przybylo et al. (2022). PCA reveals three distinct clusters, while UMAP forms more fragmented groupings. This aligns with the strong performance of the logistic regression on CPI-21K and suggests our model is learning approximately linear separable morphological features without explicit guidance. For (2), we trained a baseline classifier using the CPI-21K dataset, using 13 extracted geometric features as predictors. These geometric features are derived using traditional image processing techniques, and include features such as aspect ratio, laplacian blur, and circularity, among others. Further details about these features can be found in Przybylo et al. (2022). The feature-based logistic regression performed with a top-1 accuracy of 65%, which is much lower than the 84% accuracy from our logistic regression validation using our best model (see Table 1).
@@ 3.2 Application: quantifying ice crystal diversity
After validating the learned representations, we applied them to a downstream science task: quantifying ice habit diversity in real-world clouds. Existing methods to quantify habit diversity rely on pre-designated classes and assumptions about certain morphological features, whereas SSL-driven embeddings enable a purely data-driven approach without any prior assumptions. Since our representations follow vMF distributions, the most appropriate metric to characterize diversity is the $\kappa$ (i.e., “concentration") metric (Sra, 2012) (details in Appendix A.3). Using CPI-ENV-500K, we analyzed how $\kappa$ varies as a function of air temperature, particle size, and campaign. Figure 3a generally shows increasing habit diversity with increasing temperatures, and Figure 3b shows decreasing diversity with increasing particle size. Additionally, we see a wide spread between campaigns, highlighting the variability in crystal diversity between different cloud systems (see Appendix A.2).
@@ 4 Conclusion
We used the iBOT-vMF SSL vision transformer (Govindarajan et al., 2023) to learn ice crystal representations from a large CPI dataset. Standard SSL pre-training can be computationally expensive. Through data curation and by leveraging ImageNet pre-trained weights, we outline an efficient pre-training pipeline. This resulted in robust embeddings with strong linear predictive power for downstream habit classification, validating our learned representations. These latent representations enable fully data-driven pipelines to characterize ice crystal morphology, reducing dependence on pre-defined classes and expert-driven assumptions. As a case study, we revealed systematic ice crystal diversity variations with temperature, particle size, and campaign, and quantified intra- and inter-class similarity across habit types. For future work, we will explore how learned embeddings can support anomaly detection, identifying mislabeled or rare habits, in addition to further linking microphysical properties to thermodynamic histories. This work demonstrates that SSL can effectively capture ice crystal morphology as latent representations, providing a scalable framework to reduce microphysical uncertainties and improve the representation of ice-containing clouds in climate models.
@@ A.3 Metrics details
For a random $p$-dimensional unit vector $\bm{x_{i}}$, the von Mises-Fisher (vMF) probability distribution is given by $f(\bm{x_{i}};\bm{\mu},\kappa)=C_{p}(\kappa)\exp(\kappa\bm{\mu}^{T}\bm{x_{i}})$, where $\bm{\mu}$ is a mean vector with $\|\bm{\mu}\|=1$, $\kappa$ is a scalar concentration parameter that measures isotropic precision, and $C_{p}(\kappa)$ is a normalizing constant. A higher value of the parameter $\kappa$ denotes a higher concentration of samples around the mean vector $\bm{\mu}$ and lower variance or diversity. On the other hand, a lower value of the parameter $\kappa$ denotes a lower concentration and consequently a higher variance or diversity. We used the following equations to estimate $\kappa$ for a set of embedding vectors [Sra, 2012]:
@@ A.5.1 Self-supervised Pre-training
We use the public codebase of iBOT  [^1] https://github.com/bytedance/ibot/ and use the vMF normalized formulation proposed in Govindarajan et al. [2023]. We modified the image augmentations used during the pre-training based on their suitability to CPI images. Since, CPI images are monochromatic, we remove the jitter to the image saturation and hue. In addition to the random horizontal flip, we added a random vertical flip with a probability of $0.5$, as the crystals can be freely rotated in space. In the existing random resized crop augmentation, we reduced the change in aspect ratio of the crop by setting the new aspect ratio to be in the range of $(0.9,1.1)$. The ice crystals contain spikes of varying thickness which is an important distinguishing feature of the crystal and we want this information to be preserved in the learned representations. We use this modified set of augmentations for all the pre-training experiments that we conducted. Other hyperparameter settings for both the standard pre-training setup and the shorter and more efficient pre-training setup are provided in Table 3. In the efficient pre-training setup, we initialize the model with the weights from a model pre-trained on ImageNet-1K dataset using the iBOT method [Zhou et al., 2022].

### v3

@@ 1 Introduction
Clouds are one of the largest sources of uncertainty in climate models [1, 2]. They are notoriously difficult to represent accurately in models, and ice-containing clouds are especially challenging due to highly diverse properties such as crystal morphology [3]. Ice microphysical properties alter particle-radiation interactions and aerodynamics at the single-particle scale; and influence global radiative forcing, precipitation, and spatiotemporal distributions of clouds through a cascade of multiscale interactions [4, 5]. Improving our understanding of clouds is crucial, since uncertainties in future cloud behavior largely drive the overall uncertainty of future climate projections [6, 7, 8].
One important way to constrain ice microphysical properties is to take in situ measurements. For example, to understand the distribution of ice crystal habit (i.e., shape), millions of images of cloud particles have been taken on numerous airborne campaigns using cloud particle imagers (CPI) [9]. Historically, image processing techniques have been used to extract microphysically-relevant properties from CPI images [10, 11], and more recently, supervised ML has been used to improve predictions of particle properties [12]. However, unsupervised ML has largely been underutilized in the context of analyzing CPI data and in situ microphysical observations more broadly.
@@ 2.2 Efficient self-supervised pre-training
Self-supervised learning (SSL) is an effective approach to learn informative representations of unlabeled data [13, 14]. Such representations enable various downstream analyses as well as efficient predictions with minimum labeling efforts [15]. CPI images exhibit natural clusters characterized by ice habits. Hence, we consider state-of-the-art, clustering-based SSL methods from the DINO family [16, 17], which are trained to assign each image to a cluster such that its augmented views, obtained by applying data augmentations, are also assigned to the same cluster. Govindarajan et al. [18] showed that the representations constitute a mixture model of von Mises-Fisher (vMF) distributions. We use the iBOT-vMF method [18] to pre-train our models and use the small Vision Transformer model architecture with a patch size of $16$ (more details in A.5.1). We evaluate the learned representations with the downstream task of classifying the CPI-21K dataset.
@@
The above experiments showed that data curation is important and that ImageNet pre-trained models work reasonably well on CPI data. Also, pre-training on well-curated datasets in the target domain demonstrated potential for improved performance with better evaluation results. With this motivation, we investigated if we could pre-train models for CPI data more efficiently. Specifically, we pre-train a model on the curated CPI-H-1M dataset using iBOT-vMF for only 10 epochs and initialize the model with the ImageNet pre-trained model weights. Given a well-curated dataset like CPI-H-1M, this approach is $\sim 30\times$ more compute efficient than directly pre-training on the large CPI-3M dataset and resulted in the best performance based on validation of learned representations (see Table 1).
Table 1: Comparison of self-supervised learning models on the task of classifying the CPI-21K dataset using kNN and logistic regression classifiers. We use ViT-Small model architecture for all results and report the Top-1 accuracy metric.
@@ 3.1 Quality of learned representations
To supplement the validation results in Table 1, here we (1) confirm if clusters match expert habit labels, and (2) compare to a feature-extraction based baseline to demonstrate the benefit of SSL representations. For (1), we used dimensionality reduction to inspect clusters in 2D space. For clarity, we used a balanced subset of 3000 CPI-3M samples, and three of the seven classes to reduce visual clutter (see Appendix A.1). UMAP and PCA were used to project the 384-dimensional SSL embeddings into 2D. Figure 2 visualizes the projections, with points colored by labels predicted by a CNN from Przybylo et al. [12]. PCA reveals three distinct clusters, while UMAP forms more fragmented groupings. This aligns with the strong performance of the logistic regression on CPI-21K and suggests our model is learning approximately linearly separable morphological features without explicit guidance. For (2), we trained a baseline classifier using the CPI-21K dataset, using 13 extracted geometric features as predictors. These geometric features are derived using traditional image processing techniques, and include features such as aspect ratio, laplacian blur, and circularity, among others. Further details about these features can be found in Przybylo et al. [12]. The feature-based logistic regression performed with a top-1 accuracy of 65%, which is much lower than the 84% accuracy from our logistic regression validation using our best model (see Table 1).
@@ 3.2 Application: quantifying ice crystal diversity
After validating the learned representations, we applied them to a downstream science task: quantifying ice habit diversity in real-world clouds. Existing methods to quantify habit diversity rely on pre-designated classes and assumptions about certain morphological features, whereas SSL-driven embeddings enable a purely data-driven approach without any prior assumptions. Since our representations follow vMF distributions, the most appropriate metric to characterize diversity is the $\kappa$ (i.e., “concentration") metric [24] (details in Appendix A.3). Using CPI-ENV-500K, we analyzed how $\kappa$ varies as a function of air temperature, particle size, and campaign. Figure 3a generally shows increasing habit diversity with increasing temperatures, and Figure 3b shows decreasing diversity with increasing particle size. Additionally, we see a wide spread between campaigns, highlighting the variability in crystal diversity between different cloud systems (see Appendix A.2).
@@ 4 Conclusion
We used the iBOT-vMF SSL vision transformer [18] to learn ice crystal representations from a large CPI dataset. Standard SSL pre-training can be computationally expensive. Through data curation and by leveraging ImageNet pre-trained weights, we outline an efficient pre-training pipeline. This resulted in robust embeddings with strong linear predictive power for downstream habit classification, validating our learned representations. These latent representations enable fully data-driven pipelines to characterize ice crystal morphology, reducing dependence on pre-defined classes and expert-driven assumptions. As a case study, we revealed systematic ice crystal diversity variations with temperature, particle size, and campaign, and quantified intra- and inter-class similarity across habit types. For future work, we will explore how learned embeddings can support anomaly detection, identifying mislabeled or rare habits, in addition to further linking microphysical properties to thermodynamic histories. This work demonstrates that SSL can effectively capture ice crystal morphology as latent representations, providing a scalable framework to reduce microphysical uncertainties and improve the representation of ice-containing clouds in climate models.
Acknowledgements
We acknowledge funding from NSF through the Learning the Earth with Artificial Intelligence and Physics (LEAP) Science and Technology Center (STC) (Award #2019625). This research was also financially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP), and the Excellence Center at Linköping–Lund in Information Technology (ELLIIT). The collaboration was initiated during the ELLIIT Focus Period on Machine Learning for Climate Science in Linköping, 2024, and we thank ELLIIT for the support during the program. Computations were enabled by the Berzelius resource at the National Supercomputer Centre, provided by the Knut and Alice Wallenberg Foundation.
@@ A.3 Metrics details
For a random $p$-dimensional unit vector $\bm{x_{i}}$, the von Mises-Fisher (vMF) probability distribution is given by $f(\bm{x_{i}};\bm{\mu},\kappa)=C_{p}(\kappa)\exp(\kappa\bm{\mu}^{T}\bm{x_{i}})$, where $\bm{\mu}$ is a mean vector with $\|\bm{\mu}\|=1$, $\kappa$ is a scalar concentration parameter that measures isotropic precision, and $C_{p}(\kappa)$ is a normalizing constant. A higher value of the parameter $\kappa$ denotes a higher concentration of samples around the mean vector $\bm{\mu}$ and lower variance or diversity. On the other hand, a lower value of the parameter $\kappa$ denotes a lower concentration and consequently a higher variance or diversity. We used the following equations to estimate $\kappa$ for a set of embedding vectors [24]:
@@ A.5.1 Self-supervised Pre-training
Method overview:
We use a clustering-based self-supervised pre-training method from the DINO family, known as iBOT [17]. This uses a teacher-student self-distillation setup where both the teacher and the student have the same model architecture and are initialized with random weights. Given images $\bm{x}$} from an unlabeled image dataset, we obtain randomly augmented views of the image, ${\bm{x}}_{s}=A_{s}({\bm{x}})$ and ${\bm{x}}_{t}=A_{t}({\bm{x}})$. Here, $A_{s}$ and $A_{t}$ are random augmentations specific to the student and teacher models. The student and teacher networks consist of a backbone model (typically a Vision Transformer) and a prediction head. The backbone model is used for different downstream tasks but the prediction head (typically an MLP) is only used during pre-training and then discarded. The student and teacher models output probability distributions over $K$ pseudo-classes or clusters. The student model is trained to match the cluster assignment of the teacher, using a cross-entropy loss between the student and teacher outputs. The student model weights are updated using the loss gradients and the teacher model is updated using an exponential moving average of the student model weights. For a more detailed description and motivation behind this training methodology, we refer the reader to the original papers of DINO [16], iBOT [17] and DINO-vMF [18].
Implementation:
We use the public codebase of iBOT  [^1] https://github.com/bytedance/ibot/ and use the vMF normalized formulation proposed in Govindarajan et al. [18]. We modified the image augmentations used during the pre-training based on their suitability to CPI images. Since, CPI images are monochromatic, we remove the jitter to the image saturation and hue. In addition to the random horizontal flip, we added a random vertical flip with a probability of $0.5$, as the crystals can be freely rotated in space. In the existing random resized crop augmentation, we reduced the change in aspect ratio of the crop by setting the new aspect ratio to be in the range of $(0.9,1.1)$. The ice crystals contain spikes of varying thickness which is an important distinguishing feature of the crystal and we want this information to be preserved in the learned representations. We use this modified set of augmentations for all the pre-training experiments that we conducted. Other hyperparameter settings for both the standard pre-training setup and the shorter and more efficient pre-training setup are provided in Table 3. In the efficient pre-training setup, we initialize the model with the weights from a model pre-trained on ImageNet-1K dataset using the iBOT method [17].